OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = (x-7)/(x+3) g(x) = (-3x-7)/(x-1) So: g(f(x)) = -3(x-7/x+3)-7/(x-7/x+3)-1 f(g(x)) = (-3x-7/x-1)-7/(-3x-7/x-1)+3 But what now? I know it's something obvious, I just can't find a way to simplify.

OpenStudy (zzr0ck3r):

simplify

OpenStudy (anonymous):

How? The only thing I can find is: g(f(x)) = (-3x+21)/x+3)-7/(x-7/x+3)-1

OpenStudy (zzr0ck3r):

-3(x-7/x+3)-7/(x-7/x+3)-1 \[\frac{-3\frac{x-7}{x+3}-7}{\frac{x-7}{x+3}-1}=\] \[\frac{\frac{-3x+21-7x-21}{x+3}}{\frac{x-7-x-3}{x+3}}=\] \[\frac{-10x}{-10}=x\]

OpenStudy (zzr0ck3r):

@flixoe make sense?

OpenStudy (anonymous):

Yes that makes sense. I didn't think of creating common denominators within both the numerator and denominator. Thank you!

OpenStudy (zzr0ck3r):

yeah and then \[\frac{\frac{a}{b}}{\frac{c}{b}}=\frac{a}{c}\]